OpenStudy (anonymous):

What is the length of the missing side? A right triangle is shown. One leg is labeled nine. The hypotenuse is labeled fifteen. The second leg is labeled with a question mark. 6 twelve thirty-six square root of three hundred six

OpenStudy (anonymous):

twelve

OpenStudy (anonymous):

Use Pythagoras theorem

OpenStudy (anonymous):

Pythagorean therom=a^2+b^2=c^2 one leg is labeled, we can call it a, and hypotenuse is labeled, which is c. apply them to formula, 9^2+b^2=15^2 81+b^2=225 isolate the variable, b^2=174 to get ride of square in b, you must square root both sides. b=sqrt(174) b=13.190905...etc or just leave it as b=sqrt(174)

OpenStudy (anonymous):

ok I understand thank you

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

@Ottomaticc 225-81 = 144 !!

OpenStudy (anonymous):

opps lol calc error^ thanks

OpenStudy (anonymous):

think i put 255 by accident in my calc haha